OpenStudy (anonymous):

I have no idea how to do this problem can you guys shed some light? Sqrt[7]^2 == r^2 + r^2 - 2 * r * r * Cos[2 * Pi / 3]

OpenStudy (anonymous):

i think this is easier look at the problem:(7√)2=r2+r2−2rrcos(2π3)

OpenStudy (anonymous):

r^2(1-2cos[2pi/3] = sqrt 49 r * sqrt(1-2cos[2pi/3] ) = 49 r = 49/sqrt(1-2cos[2pi/3] ) r = 49/sqrt(2)

OpenStudy (anonymous):

o wait yeah can't really tell from how you wrote it you should put these in the equation editor

OpenStudy (anonymous):

same concept tho... not sqrt 49... would come out to just 7 instead